OpenStudy (anonymous):

Air is being pumped into a spherical balloon so that its volume increases at a rate of 100 cm^3/s. How fast is the radius of the balloon increasing when the diameter is 50cm? formula V=4/3(PI)r^3

OpenStudy (bahrom7893):

\[V=(4/3)*\pi*r^3\] \[dV/dt=4\pi*r^2(dr/dt)\] \[r=d/2=50/2=25\] \[100=4\pi*25(dr/dt)\] \[100=100*\pi(dr/dt)\] \[1=\pi(dr/dt)\] \[dr/dt=1/\pi (cm)\]

OpenStudy (bahrom7893):

Woops 1/pi (cm/s)

myininaya (myininaya):

Did you square the radius?

OpenStudy (bahrom7893):

OHH I forgot.. crap sorry